OpenStudy (anonymous):

what is sin 26 degrees 20 seconds how to calculate this??

OpenStudy (freckles):

\[1 \text{ sec}=\frac{1}{60^2} \text{ deg}\]

OpenStudy (freckles):

so \[20 \text{ sec }=\frac{20}{60^2} \text{ deg }\]

OpenStudy (freckles):

you can use your calculator to find sin(26+20/60^2 ) make sure you calculator is in deg mod

OpenStudy (freckles):

and for future reference: \[1 \text{ min }=\frac{1}{60} \text{ deg}\]

OpenStudy (freckles):

we actually call these arc seconds and arc minutes or I do I think those are the universal terms

OpenStudy (e.mccormick):

In case you are not sure why freckles did \(60^2\) there: 60 sec = 1 min 60 min = 1 degree So in degrees, 1 second is one 60th of one 60th, or \(\dfrac{1}{60^2}\) of a degree.